OpenStudy (anonymous):

use substitution method 7x+9y=21 and x=57-9y (ordered pair?)

OpenStudy (anonymous):

7(57-9y)+9y=21 399-63y+9y=21 -54y=-378 y=7 then plug in y value into either equation to find x and you'll get your ordered pair

OpenStudy (anonymous):

ordered pair (-6,7) x=-6 y=7